what happens to the volume of a fixed amount of gas if pressure ad temperature are doubled
Answer #1

Given,

P1 = P

V1 = V

T1 = T

P2 = 2 P

V2 = ?

T2 = 2 T

Formula,

P1 * V1 / T1 = P2 * V2 / T2

P * V / T = 2 P * V2 / 2 T

V2 = Final volume = V

SO, there would no change if both pressure and temperature doubled simultaneously.